Please need help right away

One container contains one blue ball and four red balls. A second container contains three blue balls and two red balls. An experiment is performed in which one of the two containers is chosen at random and then two ball are randomly chosen from it, one after the other without replacement.

a) What is the total number of outcomes of this experiment?

A 25, B 19, C 10, or D 24

b) What is the probability that two red balls are chosen?

A 6/25, B 1/4, C 6/19, or D 3/5
Answer #1

First container contains one blue Ball and four red balls

Second container contains three blue balls and two red balls

Total number of outcomes of this experiment

=( Two balls chosen from first container) or( two balls chosen from  

secondcontainer)

1)From first container chosen two balls

= (One blue Ball and one red ball) or (both red balls)

= (1 C1 * 4C1)+ (4C2)

= 4 + 6

=10

2) From second container chosen two balls

= (Two blue balls) or ( one blue and one red ball) or (two red balls)

=(3C2) + ( 3C1 * 2C1) + (2C2)

= 3+6+1

=10

Total number of outcomes of this experiment

=( Two balls chosen from first container) or( two balls chosen from  

secondcontainer)

= 10+10

=20

B)

Probability that two red balls are chosen

=( Probability of selecting first container * probability of selecting two balls from first container) + (probability of selecting second container * probablity of selecting two red balls from second container)

=(0.5 * (4C2 / 5C2) ) + ( 0.5 * ( 2C2 / 5C2))

= 6/20 + 1/20

= 7/20
Answer #2

37. A box  whole, twored balls, if two balls are selected at random find the probability that

a) one red and one white ball are selected 

b) two of the same color is selected 

c) no white balls are selected  

d) no read balls are selected